OpenStudy (anonymous):

If two distinct lines in a system of linear equations are parallel, they _______ . A. have the same slope and an infinite number of solutions B. have different slopes with one solution C. have the same slope and no solution D. have different slopes with no solution

OpenStudy (lieutenantgeneral):

A. have the same slope and an infinite number of solutions

OpenStudy (anonymous):

Can you help me with another one? @LieutenantGeneral

OpenStudy (lieutenantgeneral):

Yea.

OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using substitution: y = 4x - 10 and 5 - x = y A. Multiply the second equation by 4 B. Multiply the first equation by 2 C. Replace the left side of the first equation with the second equation D. Write the first equation in terms of x

OpenStudy (lieutenantgeneral):

D. Write the first equation in terms of x - x = 3 y = 2 B. Multiply the first equation by 2 - 4x = 12 = 4 * 3 A. Multiply the second equation by 4 - 5 * 2 = 10 C. Replace the left side of the first equation with the second equation - 5 - x = y and y = 4x - 10

OpenStudy (anonymous):

c?

OpenStudy (lieutenantgeneral):

Find this: D., B., A., & C.